[Ncep.nhc.nco_contacts] TSB Morning Rounds - Fri Mar 31, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Fri Mar 31 15:13:18 UTC 2017


Here is today's summary of NHC computer operations:
--- Mark

----------

TAFB:
The migration to RHEL6 on the NAWIPS systems continues today. NCO will be  
working on grunt and shiner at the Atlantic desk. They are also making some  
adjustments to ballyhoo to address problems that arose due to the  
conversion of the Pacific desk yesterday.

By the end of the day, the TAFB Atlantic, Pacific and surface desks will  
all be migrated. The sheephead system at the backup desk and debout at the  
standing desk might also be migrated later today if time permits.

For the weekend, keep in mind that the HSU backup desk is available to TAFB  
if unexpected problems come up on the TAFB desks.


HSU:

Other items:
The GOES-16 data will not be available over the weekend due to  
modifications in College Park. We were not able to get the dataflow from  
our local dishes set up in time to avoid the outage.

NCO plans to implement a fix to the problem that has been causing the GFS  
to come in slowly on Tuesday, April 4th.
NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170331/243ed6c8/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list